OpenStudy (rizags):

PLZ HELP!! Given: 2f(x) + f(-x) = 3x, Find: f(2)

OpenStudy (freckles):

did you try to replace the x's with 2's?

OpenStudy (freckles):

also do you know if f is even or odd?

OpenStudy (rizags):

f is a function

OpenStudy (freckles):

right do you know anything about its symmetry?

OpenStudy (freckles):

like is it even or odd?

OpenStudy (rizags):

nope

OpenStudy (freckles):

well try pluggin x=2 and also x=-2 giving you two equations

OpenStudy (freckles):

looks like you have a system of equations to solve

OpenStudy (rizags):

how does that give a system? i dont really know how to deal with functions sorry

OpenStudy (freckles):

plug in x=2: 2f(2)+f(-2)=6 plug in x=-2: 2f(-2)+f(2)=-6

OpenStudy (freckles):

We can solve for both f(2) and f(-2) but we really only need f(2) But we can solve for both for fun.

OpenStudy (freckles):

If f(2) and f(-2) look confusing to you... replace f(2) with a and f(-2) with b -- 2a+b=6 2b+a=-6

OpenStudy (freckles):

That should look a bit more familiar. That is a system of linear equations

OpenStudy (rizags):

thanks, umm, b=2 a=2??????? no way

OpenStudy (freckles):

That doesn't give you that...

OpenStudy (rizags):

ok wait i redid it SORRY

OpenStudy (rizags):

b= -6

OpenStudy (freckles):

looks good and b what f(-2)

OpenStudy (freckles):

so you found f(-2) (and i'm trying to do this fighting a kitty while it bites me)

OpenStudy (freckles):

was f(-2)* now we need our main objective which was finding a=f(2)

OpenStudy (rizags):

a=6

OpenStudy (rizags):

f(2) = 6?

OpenStudy (freckles):

looks adorable.

OpenStudy (rizags):

thanks. but how can i find the original f(x)?

OpenStudy (rizags):

becuase f(2) =6 does not imply that f(x) = 3x, right?

OpenStudy (freckles):

f(x) could be 3x because 2(3x)+(-3x)=6x-3x=3x

OpenStudy (rizags):

is there a way to find out for certain what the function is, using the given information?

OpenStudy (freckles):

well f(x)=3x is one possible answer that satisfies 2f(x)+f(-x)=3x

OpenStudy (rizags):

right, but are there (can there be) others?

OpenStudy (freckles):

Maybe

OpenStudy (rizags):

what about \[f(x) = \left| x \right|\]

OpenStudy (rizags):

that works

OpenStudy (mathmath333):

f(2)=18 is other possiblity for even function here

OpenStudy (rizags):

how did you get that?

OpenStudy (zarkon):

solve \[2f(x)+f(-1)=3x\] \[2f(-x)+f(x)=-3x\] for \(f(x)\)

OpenStudy (zarkon):

typo 2f(x)+f(-x)=3x

OpenStudy (freckles):

oh yeah that looks gorgeous zarkon

OpenStudy (mathmath333):

opps i think i read it wrong since 3x is already given

OpenStudy (rizags):

how to solve that? very new w functions

OpenStudy (zarkon):

multiply the first equation by -2 then add the equations

OpenStudy (freckles):

again rizags if it looks ugly like that you could again replace f(x) with a and replace f(-x) with b

OpenStudy (rizags):

is it -f(-x) = 2f(x)?

OpenStudy (rizags):

nope, thats wrong

OpenStudy (rizags):

what exactly should i try to eliminate when solving?

OpenStudy (freckles):

2a+b=3x a+2b=-3x --- a+2b=-3x -2(2a+b+3x) ------------ -3a+0=-9x -3a=-9x 3a=9x a=9x/3=3x So I guess f(x)=3x

OpenStudy (zarkon):

\[-2[2f(x)+f(-x)=3x]\Rightarrow-4f(x)-2f(-x)=-6x\] now add \[-4f(x)-2f(-x)=-6x\] and \[2f(-x)+f(x)=-3x\]

OpenStudy (rizags):

-3 f(x) = -9x, 3f(x) = 9x, f(x) = 3x

OpenStudy (rizags):

waiiit, then absolute value of x doesnt work?

OpenStudy (zarkon):

|x| does not work

OpenStudy (rizags):

but it appears to because 2(2) + 2 = 3(2)

OpenStudy (rizags):

where x = 2

OpenStudy (freckles):

|x|=x if x>0 |x|=-x if x<0 if x>0 then we have 2x+-x=x not 3x if x<0 then we have -2x+x=-x not 3x

OpenStudy (zarkon):

let x=-2

OpenStudy (rizags):

ohhhhHhhhhhhhhhhh got it

OpenStudy (zarkon):

if x>0 then it is true \[2f(x)+f(-x)=2|x|+|-x|=2x+x=3x\]

OpenStudy (rizags):

wait, actually, if x = -2 then 2(2) + 2 is once again 3(2)

OpenStudy (zarkon):

but you would want 3(-2)=-6

OpenStudy (rizags):

my mistake, i see now, thank you

OpenStudy (freckles):

Yes you're right @Zarkon

OpenStudy (rizags):

thanks sooo much

OpenStudy (freckles):

|-x|=x since x>0

OpenStudy (zarkon):

yep

OpenStudy (rizags):

yep got it!

OpenStudy (freckles):

or i could have said |-x|=|-1*x|=|-1|*|x|=1*|x|=1|x|=x

OpenStudy (freckles):

where x>0

OpenStudy (freckles):

just wanted to by where x>0 so people know what I'm talking about :p

OpenStudy (rizags):

yep gotcha

OpenStudy (rizags):

thanks again!